Two less than one-sixth of a certain number is 1 more than one-fourth that number. What is the number?
Answer by fractalier(6550)

(1/6)x - 2 = (1/4)x + 1
Multiply by 12 to clear fractions and get
2x - 24 = 3x + 12
and x = -36